OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. x + y = −3 y = 2x + 2 five over 3 comma 4 over 3 negative 5 over 3 comma negative 4 over 3 negative 3 over 5 comma negative 3 over 4 3 over 4 comma 3 over 5

OpenStudy (freckles):

hello have you tried subbing the second equation into the first

OpenStudy (freckles):

that is replace the y in x+y=-3 with y=2x+2 so you have x+(2x+2)=-3 can you solve this for x?

OpenStudy (anonymous):

idk how to do this

Mehek (mehek14):

x + y = −3 y = 2x + 2 if you put 2x + 2 for y in the first equation, it becomes \(x+2x+2=-3\) x + 2x =?

OpenStudy (anonymous):

idk?

Mehek (mehek14):

1 + 2?

OpenStudy (anonymous):

3

Mehek (mehek14):

so 3x x + 2x = 3x

Mehek (mehek14):

\(3x + 2 = -3\)

Mehek (mehek14):

-3 - 2?

OpenStudy (anonymous):

-3-2=-5